In a certain gas system, the pressure remains constant at 150 kPa even though 1000 J of heat is added. The volume change is 1.8 L. How much does the internal energy in joules of the system change.
1
Expert's answer
2020-11-09T10:20:56-0500

The heat added to the system:

dQ = dU + pdV = dU + p(V2 – V1)

dU – change in internal energy

p – preassure

V1 – initial volume

V2 – final volume

"dU = dQ - p(V_2 \u2013 V_1)"

"dU = 1000 – (150 \times 10^3 \times 1.8 \times 10^{-3}) = 1000 – 270 = 730 \;J"

Answer: 730 J